# A Comprehensive Evaluation of Parameter-Efficient Fine-Tuning on Code Smell Detection

### v2

@@ 2.2. Full Fine-tuning
Language Models (LMs) which are pre-trained on massive general-purpose corpus via self-supervised method, can be applied to various downstream NLP tasks by adapting their rich possession of learned knowledge through full fine-tuning (qiu2020pre, ). Full fine-tuning involves updating all parameters of an LM, including multi-head attention (MHA), feed-forward neural network (FFN), layer normalization, and so on. However, as LMs scale up, full fine-tuning a large-sized LM, i.e., LLM, can be excessively costly and ultimately make it practically unfeasible (hu2024survey, ; ding2023parameter, ).
2.3. Parameter-Efficient Fine-Tuning (PEFT) Methods
@@
Instead of full fine-tuning, pre-trained LMs could be tailored to downstream tasks by Parameter-Efficient Fine-Tuning (PEFT) methods. PEFT methods involve optimizing only a subset of model parameters or incorporating external modules for new tasks while keeping the majority of parameters frozen. Compared to full fine-tuning, PEFT methods can reduce the substantial computational and data demands of LLMs and achieve comparable or even better results (han2024parameter, ).
@@
The right side of Figure 2 illustrates several state-of-the-art PEFT methods employed in this study, including prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$. These methods were selected based on several considerations. First, they are all supported by the widely used Hugging Face PEFT library [^1] https://github.com/huggingface/peft, which provides a standardized and efficient framework for applying PEFT to LMs. Second, according to the PEFT library’s documentation [^2] https://huggingface.co/docs/peft/index, PEFT methods are categorized into three main types: Prompt-based methods, LoRA methods, and $(IA)^{3}$. The four selected methods span these categories, making them representative of the major approaches currently available. Finally, these methods are not only recent but also widely adopted in contemporary research (weyssow2023exploring, ; li2024comprehensive, ; mannisto2025comparative, ; zhou2024empirical, ; taylor2022galactica, ; esmaeili2024empirical, ; wistuba2024choice, ). Their demonstrated effectiveness across various tasks and their popularity in both the NLP and SE communities make them ideal candidates for this study. By selecting these four methods, we ensure that our experiments cover a range of the most promising and commonly adopted techniques in PEFT.
@@
Prompt tuning (lester2021power, ) directly appends learnable vectors, known as soft prompts, to the initial input embedding layer. When fine-tuning PLMs with prompt tuning, only the soft prompts are updated, while the remaining model weights and architecture are kept fixed (not updated).
@@
Prefix tuning (li2021prefix, ) introduces adjustable prefix (i.e., a sequence of task-specific vectors) across all Transformer layers. Specifically, these trainable prefix vectors are prepended to the key $K$ and value $V$ of all the MHA mechanisms. Prefix tuning adopts a reparameterization strategy, utilizing a multilayer perceptron (MLP) layer to generate the added vectors instead of optimizing them in a straightforward manner (han2024parameter, ).
@@
LoRA (Low-Rank Adaptation) (hu2022lora, ) injects trainable low-rank decomposition matrices into the query $W_{q}$, key $W_{k}$, and value $W_{v}$ projection matrices of the MHA. For a pre-trained weight matrix $W_{0}\in\{W_{q},W_{k},W_{v}\}$, an additive operation $W_{0}+\Delta W=W_{0}+BA$ is performed, where the rank of matrices $B$ and $A$ is considerably smaller than that of $W_{0}$, consequently effectively reducing the number of trainable parameters while preserving the essential characteristics of the original matrices.
@@
$\mathbf{(IA)^{3}}$ (liu2022few, ) (Infused Adapter by Inhibiting and Amplifying Inner Activation) incorporates three learnable rescaling vectors, $l_{k}$, $l_{v}$, and $l_{ff}$, to rescale the key $K$ and value $V$ in MHA and the inner activations in FFN. The operations conducted within the MHA can be outlined as follows:
[…]
@@
3. Methodology
3.1. Research Goal
Figure 2 illustrates the architecture of our LLM-based code smell detector using PEFT methods. We frame the code smell detection task as a binary classification problem, where the input is a piece of code. A sequence classification head is added on top of the LLM to handle the classification task. For model fine-tuning, we use four distinct PEFT methods. The model then outputs either 00 or $1$, representing the absence or presence of the specific type of code smell.
3.2. Research Questions
RQ1: What is the effectiveness of different PEFT methods and full fine-tuning on various SLMs and LLMs? Previous studies have shown remarkable results using PEFT methods to fine-tune SLMs for software engineering tasks like code clone detection and defect detection (liu2023empirical, ; liu2024delving, ; li2024comprehensive, ; wang2023one, ; weyssow2023exploring, ). However, these studies have not paid attention to the effectiveness of PEFT methods for code smell detection. With this RQ, we aim to compare and validate the effectiveness of PEFT methods across multiple LMs of varying parameter sizes (see Section 3.4) specifically for code smell detection.
RQ2: What is the effectiveness of PEFT methods with different hyper-parameters? When fine-tuning LMs for downstream tasks, the configuration of hyper-parameters plays a crucial role in determining the final results by affecting the number of trainable parameters. In prompt tuning and prefix tuning, the number of virtual tokens serves as a critical hyper-parameter (lester2021power, ; li2021prefix, ). For LoRA, the key hyper-parameter is the rank $r$ (hu2022lora, ), while in $(IA)^{3}$, it is the selection of activated modules (liu2022few, ). Due to our limited computational resources and the popularity of LoRA compared to other PEFT methods (gupta2024beyond, ; zhang2024star, ), we decided to employ LoRA to fine-tune a SLM and an LLM with different $r$ values in a range of {8, 32, 64, 640, 1,280, 2,560} to explore the effect of hyper-parameter settings on PEFT methods for code smell detection.
RQ3: What is the effectiveness of PEFT methods in low-resource scenarios? Low-resource scenarios, characterized by constraints on the amount of training data, pose significant challenges to fine-tuning LMs for specific tasks. Due to the scarcity of real-world datasets (zakeri2023systematic, ), these scenarios are common in tasks such as code smell detection. Therefore, it is crucial to investigate how the number of training samples influences the effectiveness of PEFT methods for code smell detection.
RQ4: What is the effectiveness of PEFT methods compared to state-of-the-art code smell detection approaches? Existing code smell detectors include traditional heuristic-based and DL-based methods. Additionally, the use of general-purpose LLMs with In-Context Learning (ICL) has become increasingly popular in SE tasks (zhou2024large, ; shin2023prompt, ; li2025large, ; xu2024unilog, ), and such models can also be leveraged for code smell detection. With this RQ, we aim to assess how LMs utilizing PEFT methods compare against heuristic-based code smell detectors, established DL techniques, and general-purpose LLMs under zero- and few-shot settings for the task of code smell detection.
3.3. Dataset Construction
In this section, we elaborate on the process with the following steps of constructing the high-quality and balanced datasets for four types of code smells used to fine-tune both small and large LMs for code smell detection. We created our own datasets because most existing datasets only offer code metrics related to the code smells instead of code snippets, which are not usable for our purposes (sharma2021qscored, ; binh2022mlcodesmell, ; madeyski2023detecting, ). While some code smell datasets provide valid code samples, they do not cover the specific types of code smells we focus on (e.g., Complex Conditional and Complex Method) (madeyski2020mlcq, ). To better align with current software development practices, we constructed new datasets based on trending GitHub projects created within the past six years. Furthermore, since data imbalance may negatively impact model performance on code smell detection (khleel2024improving, ; nandini2025improving, ; fontana2016comparing, ; nunes2024tuning, ), we ensured that our constructed datasets contain an equal number of positive and negative samples. Figure 3 illustrates the overview of the dataset construction process.
@@ 3.3. Dataset Construction
3.3.1. Selected Code Smells
@@
We selected CC, CM, FE, and DC because they are among the most extensively studied code smells in prior research on ML- and DL-based code smell detection (e.g., (sharma2021code, ; ho2023fusion, ; zhang2024code, ; alazba2023deep, ; fontana2016comparing, ; barbez2020machine, ; kim2017finding, ; liu2023deep, ; liu2018deep, ; skipina2024automatic, ; yang2024feature, ; fontana2013code, ; madeyski2020mlcq, )). Moreover, these code smells are frequently observed in real-world software projects and are known to significantly impact code quality, maintainability, and the likelihood of introducing defects (fowler1999refactoring, ). For example, CC and CM can lead to convoluted control flows that hinder code readability and increase cognitive load on developers, while FE and DC often indicate bad modularization and poor class design, which complicate maintenance and evolution (fowler1999refactoring, ; sjoberg2012questioning, ; khomh2009exploratory, ).
Another reason for choosing these specific smell types is the availability of detection tools, such as DesigniteJava (designitejava, ) and PMD (pmd, ), which facilitates the collection of candidate code samples potentially exhibiting these smells. We excluded overly simple smells (e.g., Long Method) that can be easily identified using basic heuristics. In contrast, detecting the four selected code smells requires LLMs to capture subtle features, consider aspects such as cohesion and coupling, and analyze the overall structure of a code sample (sharma2021code, ), which expands the ambition and enhances the applicability of this study.
At the same time, identifying these selected code smells allows us to assess the effectiveness of LMs fine-tuned with PEFT methods in both method-level and class-level code understanding. CC, CM, and FE are generally considered method-level code smells, while DC represents a class-level code smell. However, detecting FE requires comprehension of not only the focal method but also its surrounding class context and inter-class dependencies (fowler1999refactoring, ; tsantalis2009indentification, ; liu2023deep, ; alkharabsheh2021analysing, ). Therefore, it also demands class-level code understanding. The selection of these four code smells allows our study to address a key gap in existing research related to PEFT methods (li2024comprehensive, ; liu2023empirical, ; wang2023one, ; liu2024delving, ; weyssow2023exploring, ), which has primarily concentrated on the effectiveness of PEFT techniques on method-level code-related tasks.
3.3.2. Java Repository Selection
@@
3.3.3. Potential Code Smell Detection
DesigniteJava is an open-source tool capable of detecting 18 design smells and 10 implementation smells, whereas PMD is a rule-based static analysis tool that identifies common programming issues, including code smells. Both of them have been extensively used in prior studies (sharma2021qscored, ; shah2023mining, ; barbosa2020revealing, ; fernandes2016review, ; aniche2018code, ; wu2024ismell, ). In this paper, we leveraged DesigniteJava to detect potential instances of CC, CM, and FE, and used PMD to detect potential instances of DC. The output of this step is a list of detected code smells, each annotated with the corresponding project, package, class, and, where applicable, method name. It is worth noting that the code smell detection results from DesigniteJava and PMD may not be entirely accurate, as the detection results can contain both false positives and false negatives. False positives occur when the tool incorrectly flags code as exhibiting certain type of code smell, while false negatives happen when actual code smells are missed. These inaccuracies stem from the rule-based and heuristic nature of DesigniteJava and PMD, which relies on predefined criteria that may not capture all nuances of the code.
3.3.4. Data Preparation
CodeSplitJava (codesplitjava, ) is a utility program designed to extract methods or classes from Java source code and organize them into separate files. We employed CodeSplitJava to parse the 401 Java repositories collected in Section 3.3.2, generating individual code files for each method or class. The file paths reflect the hierarchical structure of the original code (i.e., namespaces and packages are converted into nested folders). By aligning the list of detected code smells obtained in Section 3.3.3 with the paths of the code files generated by CodeSplitJava, we were able to collect code snippets that were labeled by heuristic-based detectors as either containing code smells or being free of them.
@@ 3.3.4. Data Preparation
We removed any duplicate samples within each type of code smell to ensure uniqueness. Additionally, we limited the maximum token length of the samples to 1,024, filtering out any that exceeded this limit. This decision was made to accommodate the input constraints of the models used in our experiments (see Section 3.6). Specifically, the input window size for the LMs was set to 1,024 tokens to align with GPU memory limitations. Samples exceeding 1,024 tokens undergo truncation, which may result in the loss of critical contextual information. For instance, key sections of code relevant to the detection of code smells could be omitted, potentially impacting the model’s performance. Therefore, we set a hard limit of 1,024 tokens to avoid such issues, ensuring that all samples in our dataset can be fully utilized by the models without sacrificing the integrity of their content. Table 1 provides the number of potential positive samples after each step for the four types of code smells. Note that although FE is a method-level code smell, we retained the entire class that contains the method as the sample. This decision aligns with the definition of FE (see Section 3.3.1), which emphasizes the method’s interaction with its enclosing class. Providing the full class context ensures that the code sample contains sufficient structural and semantic information for FE detection.
3.3.5. Manual Review
@@
Table 2. Number of positive and negative samples after manual review
Code Smell #Positive #Negative Total
Complex Conditional 4,604 4,604 9,208
Complex Method 5,748 5,748 11,496
Feature Envy 307 307 614
Data Class 6,922 6,922 13,844
In order to ensure the consistency of data annotation, a pilot review was performed by the first and the fourth authors before conducting the formal manual review process. For each code smell type (i.e., CC, CM, FE, and DC), 30 potential positive samples and 30 potential negative samples labeled by heuristic-based tools were randomly selected. The two authors independently examined these samples and then compared their labeling results. Cohen’s Kappa coefficient (jacob1960coefficient, ) was used to assess inter-rater agreement between the two authors. The Kappa coefficient for CC was 0.83, for CM was 0.80, for FE was 0.77, and for DC was 0.87. All values exceeded the threshold of 0.7, indicating a strong level of agreement between the annotators. Following the pilot review, the two authors held a discussion to reconcile discrepancies and formalize the labeling criteria below:
  • CC: Labeled as positive if the method contains at least one conditional statement that either (i) has a nesting depth of three or more, making the method hard to read, or (ii) includes compound boolean expressions involving three or more operands, exhibiting complex logic that makes the condition challenging to evaluate and understand.
@@
These criteria were grounded in established definitions from prior studies (sharma2017house, ; fowler1999refactoring, ; fontana2016comparing, ; pietrzak2006leveraging, ) and refined based on insights from our pilot study. During the formal manual review, the two authors collaboratively labeled all samples (i.e., each sample was assigned to one of the two authors), which took approximately 300 person-hours in total. Whenever questions or ambiguities arose, the second author, who is an expert in code smell detection, was consulted. In such cases, all three authors discussed the sample until consensus was reached, ensuring annotation reliability and consistency. The final count of positive and negative samples, following the manual review, is presented in Table 2.
3.3.6. Data Split
In this step, we split the code samples of CC, CM, FE, and DC into training, validation, and test sets following an 8:1:1 ratio, as employed in previous studies (yang2024federated, ; liu2023benchmarking, ). For each type of code smell, we ensured that the number of positive and negative samples remains equal across the training, validation, and test sets, thereby maintaining class balance throughout the entire dataset. Table 3 outlines the number of code samples for each code smell type after the data split. We then used these clean and high-quality datasets as our final datasets for a series of experiments.

### v3

@@ 2.3.1. Full Fine-Tuning
Language Models (LMs) which are pre-trained on massive general-purpose corpus via self-supervised method, can be applied to various downstream NLP tasks by adapting their rich possession of learned knowledge through full fine-tuning (qiu2020pre). Full fine-tuning involves updating all parameters of an LM, including multi-head attention (MHA), feed-forward neural network (FFN), layer normalization, and so on. However, as LMs scale up, full fine-tuning a large-sized LM, i.e., LLM, can be excessively costly and ultimately make it practically unfeasible (hu2024survey; ding2023parameter).
2.3.2. Parameter-Efficient Fine-Tuning (PEFT) Methods
@@
Instead of full fine-tuning, pre-trained LMs could be tailored to downstream tasks by Parameter-Efficient Fine-Tuning (PEFT) methods. PEFT methods involve optimizing only a subset of model parameters or incorporating external modules for new tasks while keeping the majority of parameters frozen. Compared to full fine-tuning, PEFT methods can reduce the substantial computational and data demands of LLMs and achieve comparable or even better results (han2024parameter).
In this study, we employed the Hugging Face PEFT library [^1] https://github.com/huggingface/peft to implement and evaluate all PEFT methods. This library, developed by the Hugging Face team, offers a standardized and efficient framework that supports several widely adopted PEFT techniques and has been extensively used in recent research (li2024comprehensive; weyssow2023exploring; wang2025beyond; storhaug2024parameter). In addition, it is fully compatible with the Transformers library, allowing seamless integration with open-source LMs available on the Hugging Face platform. This unified setup ensures reproducibility, simplifies implementation, and enables fair comparisons across different PEFT strategies. Therefore, in this work, we focus exclusively on the PEFT methods provided by the PEFT library. The right side of Figure 5 illustrates the four state-of-the-art PEFT techniques adopted in our study: prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$.
@@
These methods were selected based on several considerations. First, they are all natively supported by the PEFT library, allowing stable fine-tuning of LMs while avoiding out-of-memory errors. Second, according to the official documentation [^2] https://huggingface.co/docs/peft/index, PEFT methods are categorized into three main types: Prompt-based methods, LoRA methods, and $(IA)^{3}$. The four selected methods span these categories, making them representative of the major approaches currently available. Finally, these methods are not only recent but also widely adopted in contemporary research (weyssow2023exploring; li2024comprehensive; mannisto2025comparative; zhou2024empirical; taylor2022galactica; esmaeili2024empirical; wistuba2024choice). Their demonstrated effectiveness across various tasks and their popularity in both the NLP and SE communities make them ideal candidates for this study. By selecting these four methods, we ensure that our experiments cover a range of the most promising and commonly adopted techniques in PEFT.
@@
Prompt tuning (lester2021power) directly appends learnable vectors, known as soft prompts, to the initial input embedding layer. When fine-tuning PLMs with prompt tuning, only the soft prompts are updated, while the remaining model weights and architecture are kept fixed (not updated).
@@
Prefix tuning (li2021prefix) introduces adjustable prefix (i.e., a sequence of task-specific vectors) across all Transformer layers. Specifically, these trainable prefix vectors are prepended to the key $K$ and value $V$ of all the MHA mechanisms. Prefix tuning adopts a reparameterization strategy, utilizing a multilayer perceptron (MLP) layer to generate the added vectors instead of optimizing them in a straightforward manner (han2024parameter).
@@
LoRA (Low-Rank Adaptation) (hu2022lora) injects trainable low-rank decomposition matrices into the query $W_{q}$, key $W_{k}$, and value $W_{v}$ projection matrices of the MHA. For a pre-trained weight matrix $W_{0}\in\{W_{q},W_{k},W_{v}\}$, an additive operation $W_{0}+\Delta W=W_{0}+BA$ is performed, where the rank of matrices $B$ and $A$ is considerably smaller than that of $W_{0}$, consequently effectively reducing the number of trainable parameters while preserving the essential characteristics of the original matrices.
@@
$\mathbf{(IA)^{3}}$ (liu2022few) (Infused Adapter by Inhibiting and Amplifying Inner Activation) incorporates three learnable rescaling vectors, $l_{k}$, $l_{v}$, and $l_{ff}$, to rescale the key $K$ and value $V$ in MHA and the inner activations in FFN. The operations conducted within the MHA can be outlined as follows:
(1) $\text{MHA(x)}=Softmax\Biggl(\frac{Q(l_{k}\odot K^{T})}{\sqrt{d_{k}}}\Biggr)(l_{v}\odot V)$
@@
3. Dataset Construction
In this section, we elaborate on the process of constructing high-quality and balanced datasets for four types of code smells, rendering them LM-ready and tailored specifically for fine-tuning both small and large LMs.
As discussed in Section 2.2, existing code smell datasets suffer from several critical limitations, including the absence of raw source code, insufficient coverage of complex smells, and label noise. To address these issues and ensure reproducibility, we do not rely on any existing datasets. Instead, we create our benchmark from scratch by mining recent Java repositories from GitHub, providing a solid foundation for the subsequent fine-tuning experiments on code smell detection.
Our construction pipeline is guided by three key objectives: (1) Source-Code Centricity: ensuring all samples contain compilable raw code tokens; (2) Diversity: covering four non-trivial code smells (i.e., Complex Conditional, Complex Method, Feature Envy, Data Class); and (3) Reliability: incorporating a rigorous two-stage manual review to alleviate the “Oracle Problem”. Figure 3 provides an overview of the dataset construction process.
@@ 3. Dataset Construction
3.1. Selected Code Smells
@@
  • Complex Conditional (CC): This code smell occurs when a conditional statement is excessively intricate (sharma2017house).
  • Complex Method (CM): This code smell occurs when a method exhibits high cyclomatic complexity (sharma2017house).
  • Feature Envy (FE): This code smell occurs when a method is more interested in a class other than the one it actually is in (fowler1999refactoring; fontana2016comparing).
  • Data Class (DC): This code smell occurs when a class contains only fields and basic methods for accessing them (fowler1999refactoring; fontana2016comparing).
We selected CC, CM, FE, and DC because they are among the most extensively studied code smells in prior research on ML- and DL-based code smell detection (e.g., (sharma2021code; ho2023fusion; zhang2024code; alazba2023deep; fontana2016comparing; barbez2020machine; kim2017finding; liu2023deep; liu2018deep; skipina2024automatic; yang2024feature; fontana2013code; madeyski2020mlcq)). Moreover, these code smells are frequently observed in real-world software projects and are known to significantly impact code quality, maintainability, and the likelihood of introducing defects (fowler1999refactoring). For example, CC and CM can lead to convoluted control flows that hinder code readability and increase cognitive load on developers, while FE and DC often indicate bad modularization and poor class design, which complicate maintenance and evolution (fowler1999refactoring; sjoberg2012questioning; khomh2009exploratory).
Another reason for choosing these specific smell types is the availability of detection tools, such as DesigniteJava (designitejava) and PMD (pmd), which facilitates the collection of candidate code samples potentially exhibiting these smells. We excluded overly simple smells (e.g., Long Method) that can be easily identified using basic heuristics. In contrast, detecting the four selected code smells requires LLMs to capture subtle features, consider aspects such as cohesion and coupling, and analyze the overall structure of a code sample (sharma2021code), which expands the ambition and enhances the applicability of this study.
At the same time, identifying these selected code smells allows us to assess the effectiveness of LMs fine-tuned with PEFT methods in both method-level and class-level code understanding. CC, CM, and FE are generally considered method-level code smells, while DC represents a class-level code smell. However, detecting FE requires comprehension of not only the focal method but also its surrounding class context and inter-class dependencies (fowler1999refactoring; tsantalis2009indentification; liu2023deep; alkharabsheh2021analysing). Therefore, it also demands class-level code understanding. The selection of these four code smells allows our study to address a key gap in existing research related to PEFT methods (li2024comprehensive; liu2023empirical; wang2023one; liu2024delving; weyssow2023exploring), which has primarily concentrated on the effectiveness of PEFT techniques on method-level code-related tasks.
3.2. Java Repository Selection
@@
3.3. Potential Code Smell Detection
DesigniteJava (designitejava) is an open-source tool capable of detecting 18 design smells and 10 implementation smells, whereas PMD (pmd) is a rule-based static analysis tool that identifies common programming issues, including code smells. Both of them have been extensively used in prior studies (sharma2021qscored; shah2023mining; barbosa2020revealing; fernandes2016review; aniche2018code; wu2024ismell). In this paper, we leveraged DesigniteJava to detect potential instances of CC, CM, and FE, and used PMD to detect potential instances of DC. The output of this step is a list of detected code smells, each annotated with the corresponding project, package, class, and, where applicable, method name. It is worth noting that the code smell detection results from DesigniteJava and PMD may not be entirely accurate, as the detection results can contain both false positives and false negatives. False positives occur when the tool incorrectly flags code as exhibiting certain type of code smell, while false negatives happen when actual code smells are missed. These inaccuracies stem from the rule-based and heuristic nature of DesigniteJava and PMD, which relies on predefined criteria that may not capture all nuances of the code.
3.4. Data Preparation
CodeSplitJava (codesplitjava) is a utility program designed to extract methods or classes from Java source code and organize them into separate files. We employed CodeSplitJava to parse the 401 Java repositories collected in Section 3.2, generating individual code files for each method or class. The file paths reflect the hierarchical structure of the original code (i.e., namespaces and packages are converted into nested folders). By aligning the list of detected code smells obtained in Section 3.3 with the paths of the code files generated by CodeSplitJava, we were able to collect code snippets that were labeled by heuristics-based detectors as either containing code smells or being free of them.
@@ 3.4. Data Preparation
We removed any duplicate samples within each type of code smell to ensure uniqueness. Additionally, we limited the maximum token length of the samples to 1,024, filtering out any that exceeded this limit. This decision was made to accommodate the input constraints of the models used in our experiments (see Section 4.6). Specifically, the input window size for the LMs was set to 1,024 tokens to align with GPU memory limitations. Samples exceeding 1,024 tokens undergo truncation, which may result in the loss of critical contextual information. For instance, key sections of code relevant to the detection of code smells could be omitted, potentially impacting the model’s performance. Therefore, we set a hard limit of 1,024 tokens to avoid such issues, ensuring that all samples in our dataset can be fully utilized by the models without sacrificing the integrity of their content. Table 1 provides the number of potential positive samples after each step for the four types of code smells. Note that although FE is a method-level code smell, we retained the entire class that contains the method as the sample. This decision aligns with the definition of FE (see Section 3.1), which emphasizes the method’s interaction with its enclosing class. Providing the full class context ensures that the code sample contains sufficient structural and semantic information for FE detection.
3.5. Manual Review
@@
Table 2. Number of positive and negative samples at each stage of manual review
Code Smell Stage 1 Stage 2 Total
#Positive #Negative #Positive #Negative #Positive #Negative
Complex Conditional 3,715 3,186 889 1,418 4,604 4,604
Complex Method 5,699 4,120 49 1,628 5,748 5,748
Feature Envy 184 153 123 154 307 307
Data Class 5,904 1,217 1,018 5,705 6,922 6,922
In order to ensure the consistency of data annotation, a pilot review was performed by the first and the fourth authors before conducting the formal manual review process. For each code smell type (i.e., CC, CM, FE, and DC), 30 potential positive samples and 30 potential negative samples labeled by heuristics-based tools were randomly selected. The two authors independently examined these samples and then compared their labeling results. Cohen’s Kappa coefficient (jacob1960coefficient) was used to assess inter-rater agreement between the two authors. The Kappa coefficient for CC was 0.83, for CM was 0.80, for FE was 0.77, and for DC was 0.87. All values exceeded the threshold of 0.7, indicating a strong level of agreement between the annotators. Following the pilot review, the two authors held a discussion to reconcile discrepancies and formalize the labeling criteria below:
  • CC: Labeled as positive if the method contains at least one conditional statement that satisfies (i) a nesting depth of three or more or (ii) compound boolean expressions involving three or more operands, and simultaneously (iii) exhibits semantically complex logic that makes the condition difficult to evaluate and understand.
@@
These criteria were grounded in established definitions from prior studies (sharma2017house; fowler1999refactoring; fontana2016comparing; pietrzak2006leveraging) and refined based on insights from our pilot study. During the formal manual review, the two authors collaboratively labeled all samples (i.e., each sample was assigned to one of the two authors), which took approximately 300 person-hours in total. Whenever questions or ambiguities arose, the second author, who is an expert in code smell detection, was consulted. In such cases, all three authors discussed the sample until consensus was reached, ensuring annotation reliability and consistency. Table 2 summarizes the numbers of positive and negative samples collected at each stage of the manual review process.
3.6. Data Split
In this step, we split the code samples of CC, CM, FE, and DC into training, validation, and test sets following an 8:1:1 ratio, as employed in previous studies (yang2024federated; liu2023benchmarking). For each type of code smell, we ensured that the number of positive and negative samples remains equal across the training, validation, and test sets, thereby maintaining class balance throughout the entire dataset. Table 3 outlines the number of code samples for each code smell type after the data split. We then used these clean and high-quality datasets as our final datasets for a series of experiments.